The supporting cables AB and AC are oriented so that the components of the 360-lb force along AB and AC are 185 lb and 200 lb, respectively. Determine the angles α and β


1
Expert's answer
2021-06-03T12:32:13-0400


Since the cable setup is in equilibrium, the horizontal, as well as the vertical forces, should be in equilibrium.

From the given figure,

"F(AB)*sin(\alpha) = F(AC)*sin(\beta)"

Horizontal forces in equilibrium

"F(AB)*cos(\alpha) = F(AC)*cos(\beta)"

Vertical forces in equilibrium F(AB) and F(AC) are the forces in the cables AB and AC respectively. Balancing the horizontal forces,

"F = \sqrt{F_1^2+F_2^2-2F_1F_2cos\alpha}"

"cos\alpha = \large\frac{F_1^2+F_2^2-F^2}{2F_1F_2}" = "\large\frac{185^2+200^2-360^2}{2*185*200}" "=-0.74831 \to \alpha = 138^o"

"F(AB)*sin(\alpha) = F(AC)*sin(\beta)"

"\alpha + \beta = 138^o \to \alpha = 138^o-\beta"

"185cos\alpha=200cos\beta"

"185cos(138^o-\beta)=200cos\beta"

"185(cos138^ocos\beta-sin138^osin\beta)=200cos\beta"

"\beta=\large\frac{\ln \left(-\arccos \left(\frac{40\cos \left(β\right)}{37}\right)+2\pi n+β\right)}{\ln \left(138\right)}"